Q1

A mine hauls, on average, three dump truck of waste per hour to a waste dump. For a given hour, find the probability that it will haul the following number of truck loads:

a)     At most 3 trucks.

b)     At least 3 trucks.

c)      Five or more 


1
Expert's answer
2021-07-13T05:06:52-0400

Let "X=" the number of trucks per hour: "X\sim Po(\lambda)."

Given "\lambda=3."


a)

"P(X\leq 3)=P(X=0)+P(X=1)+P(X=2)"

"+P(X=3)=\dfrac{e^{-\lambda}\cdot\lambda^0}{0!}+\dfrac{e^{-\lambda}\cdot\lambda^1}{1!}"

"+\dfrac{e^{-\lambda}\cdot\lambda^2}{2!}+\dfrac{e^{-\lambda}\cdot\lambda^3}{3!}"

"=\dfrac{e^{-3}(6+18+27+27)}{6}=13e^{-3}\approx0.647232"


b)

"P(X\geq 3)=1-P(X=0)-P(X=1)-P(X=2)"

"=1-\dfrac{e^{-\lambda}\cdot\lambda^0}{0!}-\dfrac{e^{-\lambda}\cdot\lambda^1}{1!}-\dfrac{e^{-\lambda}\cdot\lambda^2}{2!}"

"=1-\dfrac{e^{-3}(2+6+9)}{2}=1-8.5e^{-3}\approx0.576810"




c)

"P(X\geq 5)=1-P(X=0)-P(X=1)-P(X=2)"

"-P(X=3)-P(X=4)=1-\dfrac{e^{-\lambda}\cdot\lambda^0}{0!}"

"-\dfrac{e^{-\lambda}\cdot\lambda^1}{1!}-\dfrac{e^{-\lambda}\cdot\lambda^2}{2!}-\dfrac{e^{-\lambda}\cdot\lambda^3}{3!}-\dfrac{e^{-\lambda}\cdot\lambda^4}{4!}"

"=1-\dfrac{e^{-3}(24+72+108+108+81)}{24}"


"=1-\dfrac{131}{8}e^{-3}\approx0.184737"